The filament of a typical incandescent lightbulb is at a temperature of about 3,000 K. For...

  1. The filament of a typical incandescent lightbulb is at a temperature of about 3,000 K. For a 100-W lightbulb, what’s the surface area (mm2) of the filament? Assume the entire 100-W goes into radiantly heating the filament (normally, it’s about 95%; the other 5% goes into producing light) and that the filament has an emissivity of one. Please explain the math
Answer #1

According to Stefans law, Q = SAT4 = vra A Q 100 - A- = отч 5,67*108 (3x10334 -4 = A= 0.2177 * 10 m 2 m2 A= 21.77 xiob L mm